OpenStudy (yrivers36):

How do I convert x^2-4x+9 to vertex form?

OpenStudy (anonymous):

Do you know how to use the quadratic formula?

OpenStudy (yrivers36):

Im not sure what u are referring too.

OpenStudy (anonymous):

Oh, sorry thought you were looking for factors. for vertex form you just complete the square. \[x^2-4x+9 = x^2 -4x + 4 + 5 = (x-2)^2 +5\]

OpenStudy (anonymous):

There are no real factors, silly!

OpenStudy (yrivers36):

Sorry not factoring. Vertex form is different